The equilibrium constant Kc for the reaction: A(g) + B(g) ⇔ 2C(g) is 0.76 at 150 °C. If 0.500 M of A is added to 0.500 M of B in a 1.00 L container, calculate the [EQ] for A, B and C
Answer #1

= H + B 0.5 0.85 +22 - -x -x 0.5-2 0.5-2 3 2x Ke= 0.76 Ki= [c]? [AI (B] is 0.76 = 12:) (0.5-x)2 . do.76 = 0.8.72 . an 0.5-X .